# Cycle products and efficient vectors in reciprocal matrices

### v1

@@ Abstract
Given an $n$-by-$n$ reciprocal (pair-wise comparison) matrix $A=[a_{ij}]$ and a positive vector $w=\left[\begin{array}[c]{ccc}w_{1}&\cdots&w_{n}\end{array}\right]^{T},$ the directed graph $G(A,w)$ has an edge from $i$ to $j$ if and only if $w_{i}\geq a_{ij}w_{j}.$ Vector $w$ is efficient for $A$ if and only if $G(A,w)$ is strongly connected if and only if (new to this literature) $G(A,w)$ contains a Hamiltonian cycle. We show an exact correspondence between such cycles and Hamiltonian cycle products in $A$ that are $\leq 1.$ This facilitates a way to describe all efficient vectors for $A$ via cycle products, and applications, such as that the set of vectors efficient for $A$ determines $A,$ and that, if $G(A,u)$ and $G(A,v)$ present a common cycle, all vectors on the line segment joining $u$ and $v$ are efficient. A sufficient condition for convexity of the efficient set of vectors follows. Order reversals in an efficient vector are also analyzed. The description of the set of efficient vectors for an $n$-by-$n$ column perturbed consistent matrix is given. It is the union of at most $(n-1)(n-2)/2$ convex subsets.
Keywords: cycle products, decision analysis, efficient vector, pair-wise comparisons, reciprocal matrix
@@ 1 Introduction
There is a graph theoretic method to decide wether $w\in\mathcal{E}(A).$ The graph $G(A,w)$ with vertex set $\{1,\ldots,n\}$ has an edge $i\rightarrow j,$ $1\leq i\neq j\leq n,$ if and only if $w_{i}\geq a_{ij}w_{j}$. First in [6], and then in [17] in a simple matricial way, it was shown that $w\in\mathcal{E}(A)$ if and only if $G(A,w)$ is strongly connected. Since $A$is reciprocal, for every pair of distinct indices $i,j\in\{1,2,\ldots,n\},$ at least one of the edges $i\rightarrow j$ or $j\rightarrow i$ is in $G(A,w).$ A directed graph with at least one of $\{i\rightarrow j,$ $j\rightarrow i\}$ as an edge (and, perhaps both) is called semi-complete [5] (if it never happens that both occur, it is called a tournament [5, 26]); so $G(A,w)$ is always semi-complete. A semi-complete digraph is strongly connected if and only if it contains a full cycle [5], i.e. a Hamiltonian cycle. We add this fact here to the graph theoretic characterization of efficiency.
@@
In [19], a way to generate all vectors in $\mathcal{E}(A)$ (inductively) was given. Here we follow a different approach to characterize $\mathcal{E}(A).$ We connect the Hamiltonian cycles in $G(A,w)$ with the Hamiltonian cycle products $\leq 1$ from matrix $A.$ By a (Hamiltonian) cycle product in $A$ we mean the product of the entries $a_{ij}$ for which $i\rightarrow j$ is an edge of the cycle. There are at most $\frac{(n-1)!}{2}$ of these products $<1$, and exactly this number if no cycle product is $1$. We show that $\mathcal{E}(A)$ is the union of at most $\frac{(n-1)!}{2}$ convex subsets. Each subset is associated with a Hamiltonian cycle product from $A<1$ (when $A$ is inconsistent). We give sufficient conditions for the convexity of $\mathcal{E}(A),$ and for the cone generated by the columns of $A$ (that is, the set of nonzero, nonnegative linear combinations of the columns of $A$) to lie in $\mathcal{E}(A).$ The former implies the latter since any column of $A$ lies in $\mathcal{E}(A)$ [18]. When the latter occurs, the efficiency of the Perron vector of $A$ and of the left singular vector of $A$ (that is, the Perron vector of $AA^{T}$) follows, as these vectors belong to the cone generated by the columns [21].
@@
Using our new description of the efficient vectors for a reciprocal matrix, we show that $A$ is determined by $\mathcal{E}(A)$. Also, we give necessary and sufficient conditions for the existence of an efficient vector associated with a Hamiltonian cycle to have no order reversals with entries along the cycle and show that, in any circumstance, there is always one efficient vector with at most one such order reversal. An efficient vector $w$ for $A\in\mathcal{PC}_{n}$ is said to exhibit an order reversal at $i,j$ if $w_{i}>w_{j}$ ($w_{i}<w_{j}$) when $a_{ij}<1$ ($a_{ij}>1$), $w_{i}=w_{j}$ when $a_{ij}\neq 1$, or $w_{i}\neq w_{j}$ when $a_{ij}=1.$ There is now a considerable literature on order reversal and vectors that minimize the number of order reversals (see, for example, [9, 13, 30]).
@@
If one off-diagonal entry of a consistent matrix, and, correspondingly, its symmetrically placed entry, is changed, so that the resulting $A$ lies in $\mathcal{PC}_{n},$ then $A$ is called a simple perturbed consistent matrix. In prior work [10], $\mathcal{E}(A)$ for any simple perturbed consistent matrix $A\in\mathcal{PC}_{n}$ was determined, and this set is defined by a finite system of linear inequalities on the entries of the vectors, implying that $\mathcal{E}(A)$ is convex for simple perturbed consistent matrices $A$. Since this class of matrices includes $\mathcal{PC}_{3},$ it follows that $\mathcal{E}(A)$ is convex for any $A\in\mathcal{PC}_{3}.$ From the inductive description of the efficient vectors for a reciprocal matrix in [19], it follows that each efficient vector for $A\in\mathcal{PC}_{4}$ is an extension of an efficient vector for one of $3$ $3$-by-$3$ principal submatrices of $A.$ Since these extensions are defined by linear inequalities in the entries of the vector, it follows that the set of efficient vectors for $A\in\mathcal{PC}_{4}$ is the union of at most $3$ convex sets. Later, in [29] the authors illustrated geometrically how $\mathcal{E}(A)$ is the union of $3$ convex sets in the $4$-by-$4$ case.
@@
In [21] we have given examples of matrices obtained from consistent matrices by changing one column (and the corresponding row reciprocally), called column perturbed consistent matrices, for which the set of efficient vectors is not convex. However, it was shown that the cone generated by the columns of such matrices $A$ is contained in $\mathcal{E}(A).$ Note that, for any $n\geq 4,$ examples of matrices $A\in\mathcal{PC}_{n}$ for which the Perron eigenvector is not efficient are known [7, 20]. In these cases, the convex hull of the columns of the matrix is not contained in $\mathcal{E}(A)$, implying that $\mathcal{E}(A)$ is not convex. Here we describe the set of efficient vectors for a column perturbed consistent matrix $A\in\mathcal{PC}_{n}$ and show that it is the union of at most $(n-1)(n-2)/2$ convex sets. The simple perturbed consistent matrices and some type of double perturbed consistent matrices are special cases of column perturbed consistent matrices. As mentioned above, the efficient vectors for the former were studied in [10]. Those for the latter were obtained in [16].
@@
[18] Suppose that $A\in\mathcal{PC}_{n}$ and $w\in\mathcal{E}(A).$ Then, if $D$ is a positive diagonal matrix ($P$ is a permutation matrix), then $Dw\in\mathcal{E}(DAD^{-1})$ ($Pw\in\mathcal{E}(PAP^{T})$).
Note that the graphs $G(A,w)$ and $G(DAD^{-1},Dw)$ coincide.
In the next section, we present a fundamental cycle theorem for efficient vectors of reciprocal matrices. In Section 3, we apply that theorem to show that $A$ may be uniquely reconstructed  from $\mathcal{E}(A)$ and in Section 4 we give some results on order reversals in an efficient vector. Then, in Section 5 we apply the fundamental cycle theorem to study the efficient vectors for a column perturbed consistent matrix.
@@ 2 The fundamental cycle theorem for efficient vectors of reciprocal matrices
Here, we present a fundamental theorem, based upon cycles, that explains $\mathcal{E}(A)$ for $A\in\mathcal{PC}_{n}.$ This permits insight into the nature of $\mathcal{E}(A),$ including a sufficient condition for convexity and an application that $A$ is determined by $\mathcal{E}(A).$
@@ 2.1 Auxiliary results
Case i) Suppose that $a_{i,i+k}<1$ and $a_{i+1,i+1+k}\leq 1.$ Let
@@
Case iib) Suppose that $k=2$, $a_{n-2,n}<1$, $a_{n-3,n-1}>1$, etc. Let
@@ 2.2 Description of the set $\mathcal{\varepsilon}_{A}(\mathcal{C})$
In what follows, given $A\in\mathcal{PC}_{n}$ and a Hamiltonian cycle $\mathcal{C}$, we denote by $\mathcal{\varepsilon}_{A}(\mathcal{C})$ the set of vectors $w$ such that $G(A,w)$ contains the cycle $\mathcal{C}$. Note that, by Theorem 1, $\varepsilon_{A}(\mathcal{C})\subseteq\mathcal{E}(A)$ and, by Lemma 4, that $\varepsilon_{A}(\mathcal{C})$ is convex. Next, we describe the sets $\mathcal{\varepsilon}_{A}(\mathcal{C}).$ We denote by $\pi_{1}(A)$ and $\pi_{<1}(A)$ the set of all the Hamiltonian cycles for $A$ with product $\leq 1$ and $<1$, respectively. Of course $\pi_{<1}(A)\subseteq\pi_{1}(A).$
@@
$\begin{array}[c]{rrr}w_{\gamma_{1}}\geq a_{\gamma_{1}\gamma_{2}}w_{\gamma_{2}},&\quad w_{\gamma_{2}}\geq a_{\gamma_{2}\gamma_{3}}w_{\gamma_{3}},&\\ \ldots,&w_{\gamma_{n-1}}\geq a_{\gamma_{n-1}\gamma_{n}}w_{\gamma_{n}},&\quad w_{\gamma_{n}}\geq a_{\gamma_{n}\gamma_{1}}w_{\gamma_{1}},\end{array}$ (2)
@@
$w_{\gamma_{1}}$ $\geq a_{\gamma_{1}\gamma_{2}}w_{\gamma_{2}}\geq\left({\displaystyle\prod\limits_{i=1}^{2}}a_{\gamma_{i}\gamma_{i+1}}\right)w_{\gamma_{3}}\geq$ (3)
@@
Suppose that $A\in\mathcal{PC}_{n}$ and that $\mathcal{C},$ as in (1), lies in $\pi_{1}(A)$. The set of positive solutions $w$ of (2), when any $n-1$ inequalities are taken to be equalities, is a singleton. Moreover, a vector in each of the $n$ singletons is extreme in $\mathcal{\varepsilon}_{A}(\mathcal{C}),$ so that $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is the cone generated by these $n$ vectors.
@@
Proof. Since a diagonal similarity does not change any cycle product in $A$, and by a diagonal similarity on $A$ any $n-1$ entries along $\mathcal{C}$ may be made $1$ (and the remaining entry is $\leq 1,$ since $\mathcal{C}\in\pi_{1}(A)$), taking into account Lemma 2, we assume that this situation occurs in order to prove the result. When the product for $\mathcal{C}$ in $A$ is $1,$ all entries in $A$ along $\mathcal{C}$ are $1$ Then, from (2) (with $\mathcal{C}$ as in (1)), $w\in\varepsilon_{A}(\mathcal{C})$ if and only, up to a positive factor, $w$ is the vector of $1$’s, implying that $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is a singleton. When the product of $\mathcal{C}$ in $A$ is $<$ $1,$ by Lemma 6, there is a $w$ satisfying (2) in which the first $n-1$ inequalities are taken to be equalities. Then, the $n$-th inequality is strict. Any vector $w^{\prime}$ obtained from $w$ by a sufficiently small decrease of $w_{\gamma_{n}},$ and agreeing with $w$ in all other components, still satisfies (2) (with $w_{\gamma_{n}}^{\prime}$ instead of $w_{\gamma_{n}}$). So $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is not a singleton. This completes the proof of the first claim.
Suppose that the product along $\mathcal{C}$ in $A$ (inconsistent) is $1,$ in which case all vectors $w$ in $\varepsilon_{A}(\mathcal{C})$ are constant. By Lemma 5, there is a Hamiltonian cycle $\mathcal{C}^{\prime}$ in $A$ such that all entries in $A$ along $\mathcal{C}^{\prime}$ are $\leq 1$ and there is at least one entry $<1.$ For $\mathcal{C}^{\prime}$ as in (1), it follows that $w$ satisfies (2), implying that $w\in\varepsilon_{A}(\mathcal{C}^{\prime}).$ When $\mathcal{C}^{\prime}\in\pi_{<1}(A),$ the last claim follows.
We observe that, if the product along a Hamiltonian cycle $\mathcal{C}$ in $A$ is $1,$ then so is the product along $\mathcal{C}^{\prime},$ and $\mathcal{\varepsilon}_{A}(\mathcal{C})=\mathcal{\varepsilon}_{A}(\mathcal{C}^{\prime}),$ in which $\mathcal{C}^{\prime}$ is the reverse cycle of $\mathcal{C}$.

### v3

@@ Abstract
We focus upon the relationship between Hamiltonian cycle products and efficient vectors for a reciprocal matrix $A$, to more deeply understand the latter. This facilitates a new description of the set of efficient vectors (as a union of convex subsets), greater understanding of convexity within this set and of order reversals in efficient vectors. A straightforward description of all efficient vectors for an $n$-by-$n$, column perturbed consistent matrix is given; it is the union of at most $(n-1)$ choose $2$ convex sets.
Keywords: cycle products, decision analysis, efficient vector, Hamiltonian cycle, pair-wise comparisons, reciprocal matrix
@@ 1 Introduction
There is a graph theoretic method to decide whether $w\in\mathcal{E}(A).$ The graph $G(A,w)$ with vertex set $\{1,\ldots,n\}$ has an edge $i\rightarrow j,$ $1\leq i\neq j\leq n,$ if and only if $w_{i}\geq a_{ij}w_{j}$. First in [6], and then in [17] in a simple matricial way, it was shown that $w\in\mathcal{E}(A)$ if and only if $G(A,w)$ is strongly connected. Since $A$is reciprocal, for every pair of distinct indices $i,j\in\{1,2,\ldots,n\},$ at least one of the edges $i\rightarrow j$ or $j\rightarrow i$ is in $G(A,w).$ A directed graph with at least one of $\{i\rightarrow j,$ $j\rightarrow i\}$ as an edge (and, perhaps both) is called semi-complete [5] (if it never happens that both occur, it is called a tournament [5, 27]); so $G(A,w)$ is always semi-complete. A semi-complete digraph is strongly connected if and only if it contains a full cycle [5], i.e. a Hamiltonian cycle. We add this fact here to the graph theoretic characterization of efficiency.
@@
In [21], a way to generate all vectors in $\mathcal{E}(A)$ (inductively) was given. Here we follow a different approach to characterize $\mathcal{E}(A).$ We connect the Hamiltonian cycles in $G(A,w)$ with the Hamiltonian cycle products $\leq 1$ from matrix $A.$ By a (Hamiltonian) cycle in $A$ we mean a sequence of $n$ entries $a_{ij}$ for which $i\rightarrow j$ are the edges of a Hamiltonian cycle; the product of these entries is the (Hamiltonian) cycle product. There are at most $\frac{(n-1)!}{2}$ of these products $<1$, and exactly this number if no cycle product is $1$. We show that $\mathcal{E}(A)$ is the union of at most $\frac{(n-1)!}{2}$ convex subsets. Each subset is associated with a Hamiltonian cycle product from $A<1$ (when $A$ is inconsistent). We give sufficient conditions for the convexity of $\mathcal{E}(A),$ and for the cone generated by the columns of $A$ (that is, the set of nonzero, linear combinations of the columns of $A$ with nonnegative coefficients) to lie in $\mathcal{E}(A).$ The former implies the latter since any column of $A$ lies in $\mathcal{E}(A)$ [18]. If the latter occurs, the efficiency of the Perron vector of $A$ and of the left singular vector of $A$ (that is, the Perron vector of $AA^{T}$) follows, as these vectors belong to the cone generated by the columns [22].
@@
Using our new description of the efficient vectors for a reciprocal matrix, we give necessary and sufficient conditions for the existence of an efficient vector associated with a Hamiltonian cycle to have no order reversals with entries along the cycle and show that, in any circumstance, there is always one efficient vector with at most one such order reversal. An efficient vector $w$ for $A\in\mathcal{PC}_{n}$ is said to exhibit an order reversal at $i,j$ if $w_{i}>w_{j}$ ($w_{i}<w_{j}$) when $a_{ij}<1$ ($a_{ij}>1$), $w_{i}=w_{j}$ when $a_{ij}\neq 1$, or $w_{i}\neq w_{j}$ when $a_{ij}=1.$ There is now a considerable literature on order reversal and vectors that minimize the number of order reversals (see, for example, [9, 13, 31]).
@@
If one off-diagonal entry of a consistent matrix, and its symmetrically placed entry, are changed, so that the resulting $A$ lies in $\mathcal{PC}_{n},$ then $A$ is called a simple perturbed consistent matrix. In prior work [10], $\mathcal{E}(A)$ for any simple perturbed consistent matrix $A\in\mathcal{PC}_{n}$ was determined, and this set is defined by a finite system of linear inequalities on the entries of the vectors, implying that $\mathcal{E}(A)$ is convex for simple perturbed consistent matrices $A$. Since this class of matrices includes $\mathcal{PC}_{3},$ it follows that $\mathcal{E}(A)$ is convex for any $A\in\mathcal{PC}_{3}.$ Recently, in [30], the authors illustrated geometrically how $\mathcal{E}(A)$ is the union of $3$ convex sets in the $4$-by-$4$ case. These facts follow from the main result in this paper (Theorem 8).
@@
In [22] we have given examples of matrices obtained from consistent matrices by changing one column (and the corresponding row reciprocally), called column perturbed consistent matrices, for which the set of efficient vectors is not convex. However, it was shown that the cone generated by the columns of such matrices $A$ is contained in $\mathcal{E}(A).$ Here we describe the set of efficient vectors for a column perturbed consistent matrix $A\in\mathcal{PC}_{n}$ and show that it is the union of at most $(n-1)(n-2)/2$ convex sets. The simple perturbed consistent matrices and some type of double perturbed consistent matrices are special cases of column perturbed consistent matrices. As mentioned above, the set of efficient vectors for the former was studied in [10] and is convex. The one for the latter was obtained in [16]. We give here an example illustrating that it may be not convex. We also present an example of a matrix $A$ that is neither a simple nor a double perturbed consistent matrix and for which $\mathcal{E}(A)$ is convex. In the more recent paper [19], we have described the efficient vectors for reciprocal matrices obtained from a consistent matrix by modifying a $3$-by-$3$ principal submatrix and have provided a class of efficient vectors when the modified block is of size greater than $3$.
@@
[18] Suppose that $A\in\mathcal{PC}_{n}$ and $w\in\mathcal{E}(A).$ Then, if $D$ is an $n$-by-$n$ positive diagonal matrix ($P$ is an $n$-by-$n$ permutation matrix), then $Dw\in\mathcal{E}(DAD^{-1})$ ($Pw\in\mathcal{E}(PAP^{T})$).
Note that the graphs $G(A,w)$ and $G(DAD^{-1},Dw)$ coincide. From the lemma we have that, if $S$ is an $n$-by-$n$ monomial matrix, then $\mathcal{E}(SAS^{-1})=S\mathcal{E}(A).$
In the next section, we present a fundamental cycle theorem for efficient vectors of reciprocal matrices. In Section 3, we apply that theorem to give some results on order reversals in an efficient vector. Then, in Section 4 we apply the fundamental cycle theorem to study the efficient vectors for a column perturbed consistent matrix. Some conclusions are presented in Section 5.
@@ 2 The fundamental cycle theorem for efficient vectors of reciprocal matrices
Here, we present a fundamental theorem, based upon cycles, that explains $\mathcal{E}(A)$ for $A\in\mathcal{PC}_{n}.$ This permits insight into the nature of $\mathcal{E}(A),$ including a sufficient condition for convexity.
@@ 2.1 Auxiliary results
Case i) Suppose that $a_{i,i+k}<1$ and $a_{i+1,i+1+k}\leq 1$ for some $i\in\{1,\ldots,n-k-1\}.$ Let
@@
Case iib) Suppose that $k=2$, $a_{n-2,n}<1$, $a_{n-3,n-1}\geq 1$, $a_{n-4,n-2}\leq 1$, $a_{n-5,n-3}\geq 1,$ etc. Let
@@ 2.2 Description of the set $\mathcal{\varepsilon}_{A}(\mathcal{C})$
In what follows, given $A\in\mathcal{PC}_{n}$ and a Hamiltonian cycle $\mathcal{C}$, we denote by $\mathcal{\varepsilon}_{A}(\mathcal{C})$ the set of vectors $w$ such that $G(A,w)$ contains the cycle $\mathcal{C}$. Note that, by Theorem 1, $\varepsilon_{A}(\mathcal{C})\subseteq\mathcal{E}(A)$ and, by Lemma 4, $\varepsilon_{A}(\mathcal{C})$ is convex. Next, we describe the sets $\mathcal{\varepsilon}_{A}(\mathcal{C}).$ We denote by $\pi_{1}(A)$ and $\pi_{<1}(A)$ the set of all Hamiltonian cycles for $A$ with product $\leq 1$ and $<1$, respectively. Of course $\pi_{<1}(A)\subseteq\pi_{1}(A).$
@@
$\begin{array}[c]{lll}w_{\gamma_{1}}\geq a_{\gamma_{1}\gamma_{2}}w_{\gamma_{2}},&\quad w_{\gamma_{2}}\geq a_{\gamma_{2}\gamma_{3}}w_{\gamma_{3}},&\ldots,\\ &&\\ \ldots,&w_{\gamma_{n-1}}\geq a_{\gamma_{n-1}\gamma_{n}}w_{\gamma_{n}},&\quad w_{\gamma_{n}}\geq a_{\gamma_{n}\gamma_{1}}w_{\gamma_{1}},\end{array}$ (2)
@@
$w_{\gamma_{1}}$ $\geq a_{\gamma_{1}\gamma_{2}}w_{\gamma_{2}}\geq\left({\displaystyle\prod\limits_{i=1}^{2}}a_{\gamma_{i}\gamma_{i+1}}\right)w_{\gamma_{3}}\geq$
@@
Suppose that $A\in\mathcal{PC}_{n}$ and that $\mathcal{C},$ as in (1), lies in $\pi_{1}(A)$. The set of positive solutions $w$ to (2), when any $n-1$ inequalities are taken to be equalities, is a singleton. Moreover, a vector in each of the $n$ singletons is extreme in $\mathcal{\varepsilon}_{A}(\mathcal{C}),$ so that $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is the cone generated by these $n$ vectors.
@@
Proof. Since a diagonal similarity does not change any cycle product in $A$, and by a diagonal similarity on $A$ any $n-1$ entries along $\mathcal{C}$ may be made $1$ (and the remaining entry is $\leq 1,$ since $\mathcal{C}\in\pi_{1}(A)$), taking into account Lemma 2, we assume that this situation occurs in order to prove the result.
When the product for $\mathcal{C}$ in $A$ is $1,$ all entries in $A$ along $\mathcal{C}$ are $1$ Then, from (2) (with $\mathcal{C}$ as in (1)), $w\in\varepsilon_{A}(\mathcal{C})$ if and only if, up to a positive factor, $w$ is the vector of $1$’s, implying that $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is a singleton. When the product of $\mathcal{C}$ in $A$ is $<$ $1,$ by Lemma 6, there is a $w$ satisfying (2) in which the first $n-1$ inequalities are taken to be equalities. Then, the $n$-th inequality is strict. Any vector $w^{\prime}$ obtained from $w$ by a sufficiently small decrease of $w_{\gamma_{n}},$ and agreeing with $w$ in all other components, still satisfies (2) (with $w_{\gamma_{n}}^{\prime}$ instead of $w_{\gamma_{n}}$). So $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is not a singleton. This completes the proof of the first claim.
Suppose that the product along $\mathcal{C}$ in $A$ (inconsistent) is $1,$ in which case all entries in $A$along $\mathcal{C}$ are $1$ (as we are assuming that there are $n-1$ such entries equal to $1$). Thus, all vectors $w$ in $\varepsilon_{A}(\mathcal{C})$ are constant. By Lemma 5, there is a Hamiltonian cycle $\mathcal{C}^{\prime}$ in $A$ such that all entries in $A$ along $\mathcal{C}^{\prime}$ are $\leq 1$ and there is at least one entry $<1.$ For $\mathcal{C}^{\prime}$ as in (1), it follows that $w$ satisfies (2), implying that $w\in\varepsilon_{A}(\mathcal{C}^{\prime}).$ Since $\mathcal{C}^{\prime}\in\pi_{<1}(A),$ the last claim follows.
We observe that, if the product along a Hamiltonian cycle $\mathcal{C}$ in $A$ is $1,$ then so is the product along the reverse cycle $\mathcal{C}^{r}$ of $\mathcal{C},$ and $\mathcal{\varepsilon}_{A}(\mathcal{C})=\mathcal{\varepsilon}_{A}(\mathcal{C}^{r})$.